# Many-Body Dissipative Particle Dynamics Simulations of Lipid Bilayers with the MDPD-MARTINI Force-Field

### v1

@@ III Results and discussion
Figure 5: Second-rank order-parameter, $P_{2}$, calculated for every bond along the lipid molecule in the case of POPC and DOPC lipids. Each bilayer consisted of 8192 lipds.
Lipid bilayers possess elastic properties, which can be expressed by the bending modulus, $k_{c}$. In particular, this reflects the intensity of undulations within the lipid membrane and can be obtained by the fluctuation spectrum [50]. By replicating the bilayers in the bilayer plane and running short simulations to reach equilibrium structures, the undulation spectrum was determined. Specifically, undulations can be described through a surface function, $u(x,y)$, which represents the average position of the two monolayers based on the positions of the phosphate head beads. Then, the undulation spectrum, $\hat{u}$, can be obtained by performing a Fourier transform on the surface function, that is:
@@
Moreover, the undulation intensities in the Fourier space are expected to follow the relation [50]
@@
where $A=L_{x}L_{y}$ is the lipid-membrane area, while $q=|\mathbf{q}|=\sqrt{q_{x}^{2}+q_{y}^{2}}$ is the wavevector length in the radial direction and $\gamma$ the surface tension. Our data, shown in Figure 4, for the DOPC and POPC bilayers exhibit the $q^{-4}$ scaling for the long wavelength modes, with a very good match between the MD and the MDPD results. Moreover, in comparison with previous results for DPPC bilayers [26], here, the $q^{-2}$ dependence is not present due to the application of the zero surface-tension condition. The $q{-4}$ dashed line was plotted by considering the experimentally obtained bending modulus $k_{c}=5^{-20}$ [51].
To further compare the MD- and MDPD-MARTINI results, the second-rank order parameter $P_{2}$ was computed, given by the following relation:
@@
IV Conclusion
The MDPD-MARTINI FF was employed for investigating structural properties of lipid bilayers with results being compared to simulations based on the widely used MD-MARTINI CG FF. The two models are in very good agreement in a range of different properties discussed in this study, thus providing further support in favor of using MDPD models, since simulations are much faster in terms of computational time and even more in terms of the simulated physical time of the systems. This particular feature opens the door for reaching larger systems and longer times scales in the simulation of complex systems, covering the full-range of chemical groups currently available in the MD-MARTINI models. This will also has the potential of accelerating the acquisition of physical-based data, which can be used in machine learning or other artificial-intelligence methods. Moreover, the FF parameters of the C3 MARTINI bead-type has been provided to the MDPD-MARTINI interaction matrix [26]. In view of the transferability of the potential based on the hydrophobic/hydrophilic characteristics of chemical groups represented by the CG beads, which is an important concept in the MARTINI coarse-graining approach and demonstrated here in the case of DPPC, POPC, and DOPC lipids. Thus, a range of different systems can be simulated by the provided set of interactions by following the MARTINI mapping of CG beads to chemical units for a range of different systems and additional ones can be investigated for further validations.
@@ Acknowledgments
Conflict of interest
There is no conflict of interest.

### v2

@@ 3 Results and Discussion
Figure 5: Second-rank order-parameter, $P_{2}$, calculated for every bond along the lipid molecule in the case of POPC and DOPC lipids. Each bilayer consisted of 8192 lipids.
Lipid bilayers possess elastic properties, which can be expressed by the bending modulus, $k_{c}$. In particular, this reflects the intensity of undulations within the lipid membrane and can be obtained by the fluctuation spectrum tarazona2013 . By replicating the bilayers in the bilayer plane and running short simulations to reach equilibrium structures, the undulation spectrum was determined. Specifically, undulations can be described through a surface function, $u(x,y)$, which represents the average position of the two monolayers based on the positions of the phosphate head beads. Then, the undulation spectrum, $\hat{u}$, can be obtained by performing a Fourier transform on the surface function, that is:
@@
Moreover, the undulation intensities in the Fourier space are expected to follow the relation tarazona2013
@@
where $A=L_{x}L_{y}$ is the lipid-membrane area, while $q=|\mathbf{q}|=\sqrt{q_{x}^{2}+q_{y}^{2}}$ is the wavevector length in the radial direction and $\gamma$ the surface tension. Our data, shown in Figure 4, for the DOPC and POPC bilayers exhibit the $q^{-4}$ scaling for the long wavelength modes, with a very good match between the MD and the MDPD results. Moreover, in comparison with previous results for DPPC bilayers Carnevale2024_MDPD_MARTINI , here, the $q^{-2}$ dependence is not present due to the application of the zero surface-tension condition. The $q^{-4}$ dashed line was plotted by considering the experimentally obtained bending modulus $k_{c}=5^{-20}$ (rawicz2000, ).
To further compare the MD- and MDPD-MARTINI results, the second-rank order parameter $P_{2}$ was computed Chremos2016 , given by the following relation:
@@
4 Conclusions
The MDPD-MARTINI-FF was employed for investigating structural properties of lipid bilayers with results being compared to simulations based on the widely used MD-MARTINI CG FF. The two models are in very good agreement in a range of different properties discussed in this study, thus providing further support in favor of using MDPD models, since simulations are much faster in terms of computational and physical times Carnevale2024_MDPD_MARTINI . This particular feature opens the door for reaching larger systems and longer times scales in the simulation of complex systems, covering the full-range of chemical groups currently available in the MD-MARTINI models. This will also have the potential of accelerating the acquisition of physical-based data, which can be used in machine learning or other artificial-intelligence methods. In various applications, the presence of soft interactions in MDPD models may pose certain limitations when hard-core interactions might be necessary. While this remains to be seen, various solutions can be envisioned in such a scenario beyond the possible use of hybrid models based on MDPD and LJ interactions. In addition, further work with MDPD models is expected to provide more information regarding any limitations of the MDPD models for specific systems, as MDPD models gain in popularity.
In this study, the FF parameters of the C3 MARTINI bead-type has been provided to the MDPD-MARTINI interaction matrix Carnevale2024_MDPD_MARTINI . In view of the transferability of the potential based on the hydrophobic/hydrophilic characteristics of chemical groups represented by the CG beads, which is an important concept in the MARTINI coarse-graining approach and demonstrated here in the case of DPPC, POPC, and DOPC lipids, a range of different systems can be simulated by the provided set of interactions. Thus, by following the MARTINI mapping of CG beads to chemical units additional systems can be investigated for further validations. Finally, methods, such as the GōMARTINI for proteinsPoma2017 , will be integrated into the general-purpose MDPD-MARTINI-FF as things progress.
@@ Acknowledgments
Data Availability Statement
Data sets generated during the current study are available from the corresponding author on reasonable request.